Taxes End of Chapter Problem 13 a. The following tax proposal has income as the tax base. Calculate the marginal tax rate for13 a. The following tax proposal has income as the tax base. Calculate the marginal tax rate for an income of $5,000 and an income of $40,000. Then, calculate the percentage of income paid in taxes for an individual with a pre‐tax income of $5,000 and for an individual with a pre tax income of $40,000. Classify the tax as being proportional, progressive, or regressive. Round answers to two places after the decimal, where applicable.

Tax Proposal: All income is taxed at 20%.

marginal tax rate; income is $5,000:

%

marginal tax rate; income is $40,000:

%

percentage paid on an income of $5,000:

%

percentage paid on an income of $40,000:

%

This tax is

proportional.

progressive.

regressive.
Answer #1

Marginal tax = 20% of $ 5000 = $ 1,000

Marginal tax = 20% of $ 40,000 = 0.2*$40,000 = $ 8,000

Marginal tax rate, when Income is $ 5,000

= (1,000/5,000)*100 = 20%

Marginal tax rate, when Income is $ 40,000

= (8,000/40,000)*100 = 20%

Percentage paid on an income of $ 5,000

=(1000/5000)*100 = 20%

Percentage paid on an income of 40,000

=(8000/40000)*1000 = 20%

The tax is Proportional

Please contact if having any query thank you. Kindly contact before rating up will be obliged to you for your generous support.